# Towards Reliable Latent Knowledge Estimation in LLMs: In-Context Learning vs. Prompting Based Factual Knowledge Extraction

### v1

@@ 4.2.2 Comparing within the same LLM family
Larger models embed more knowledge. We show in Figure 5 that, within each model family, bigger models (e.g. Llama-65B) generally outperform their smaller counterparts (e.g. Llama-13B) in terms of accuracy with an exception in the OPT family. Models within the same family are typically pre-trained on the same datasets (Biderman et al., 2023; Zhang et al., 2022; Touvron et al., 2023). Thus, this observation suggests that, when trained on identical datasets, the larger models capture a broader set of facts.
@@
Despite being trained on the same data, models might remember different facts. From these results, however, it is not clear if the larger models are subsuming smaller models in their factual knowledge, i.e., are the larger models also correct on the facts that the smaller models are correct on? To assess this, we compute the subsumption rate :
@@
i.e., the fraction of facts from known by smaller model that larger model also knows. A subsumption rate of 1 indicates that all of the smaller model’s knowledge is also contained in the larger model. To ensure a meaningful comparison across scales, we only consider models that were pre-trained using the same training data.
@@
Table 1 shows the average subsumption rate () between the largest and smallest models in a family, as well as the average accuracy, over all relations for different model families. Interestingly, is relatively low (< 0.5) for OPT, Pythia and Bloom (i.e., the larger models know less than 50% of what the smaller models know) and only reaching up to 0.8 for Gemma, Llama and Llama-2. Therefore, even though models within each family are trained on the same datasets and generally agree on the relative knowledge of different relations (Figure 6), there are differences in the knowledge of specific facts they retain from their training data.
@@
Fine-tuning reduces latent knowledge. Finally, we investigate the effects of chat-based fine-tuning on the factual knowledge of models. Base language models are often fine-tuned (using a mix of supervised and reinforcement learning Ouyang et al. (2022)) to make them better at following instructions. While prior works have shown that this makes the models better at various benchmarks, it’s unclear how such fine-tuning affects latent knowledge. Figure 7 illustrates the comparative accuracy of pre-trained models and their fine-tuned counterparts. In almost all cases, the fine-tuned models obtain lower accuracy than their base versions. This suggests that fine-tuning reduces the amount of extractable latent knowledge in the models. A similar observation was also made by Yu et al. (2024). We observe a similar trend using EIC-LKE in Appendix H.6, Figure 15. Additional results on evaluating generated outputs (using 50 tokens) in Figure 16 reveal the same pattern. To further assess if the fine-tuned models are acquiring new knowledge, we compute the subsumption rate between pre-trained and fine-tuned versions (Table 10). We find that most of the latent knowledge in fine-tuned models is already present in base models (high ), thus indicating, that fine-tuned models may not be obtaining additional knowledge.
5 Concluding Discussion
@@
In this work, we investigate a new way to estimate latent factual knowledge from an LLM. Unlike prior approaches that use prompting, our method relies on in-context learning. Our method not only addresses many reliability concerns with prompting, but it also recollects (at time significantly) more factual knowledge than prompting. In contrast to prompting, which requires relationship-specific and LLM-specific prompt engineering, our method can be applied with minimal effort to test factual knowledge of relations across a variety of structured knowledge bases and LLMs. This ability enables us to compare the latent knowledge captured by many different families of open-source LLMs; we expect our results to be of interest to designers of these LLMs. Finally, to design our in-context learning based LKE, we explore the impact of the number and ordering of correct, incorrect, and unknown examples used as inputs; our findings may be of independent interest to developing a better understanding of in-context learning.
@@
A fundamental question posed by our and prior work on estimating latent knowledge in LLMs: What does it mean for an LLM to know a fact? Suppose we tried to infer if an LLM knows the capital of Germany using the input "France Paris; Spain Madrid; Germany " and suppose the answer were Berlin. What we have learnt is that the LLM knows that the relationship between Germany and Berlin is similar to that between France and Paris or Spain and Madrid. What we have not learned is whether the LLM knows that the relation is called "capital" in English or "hauptstadt" in German. The latter is revealed by prompts such as "The capital of Germany is ". But, such prompts don’t reveal whether the LLM knows that what Berlin means to Germany is similar to what Paris means to France.
@@
Is one type of knowing facts better than other? It is difficult to answer in general. Neither type of knowing guarantees that the knowledge can be put to use in different contexts and tasks, such as when we ask the LLM where the parliament of Germany is located. Nevertheless, one clear takeaway from our study is related to how factual knowledge is latently embedded in an LLM. We show that more factual knowledge can be recollected using in-context learning, i.e., the representations of subjects and objects that share the same relationship, than by prompting with the name of their relationship.
6 Limitations
This study contributes to advancing our understanding of latent factual knowledge in LLMs through an innovative in-context learning approach. However, it is essential to acknowledge the inherent limitations of our work. While the use of in-context learning aims to mitigate the influence of prompt engineering and the reliability issues associated with previous prompting methods, it introduces its own biases based on the selection and formulation of in-context examples. We discus these in detail in Section 3. For example, the choice of which examples to include, their order, and their factual accuracy can influence model responses, and thus these in-context examples must be carefully curated for reliable latent knowledge estimation. Additionally, our study’s limitation in testing simple-format facts underlines a critical gap in assessing LLMs’ complex reasoning abilities. The knowledge estimation framework employed predominantly hinges on the LLM’s capacity to correctly recall or recognize factual information from a given set of triplets or structured prompts. This narrows the scope of evaluation to straightforward factual recall, thereby overlooking the models’ capability to engage in more sophisticated cognitive processes such as reasoning, synthesis, and inference, which we leave as open avenues for future work.
@@ Appendix A Dataset
A.1 Creation of Nobel laureates dataset from Wikidata
@@
A.2 Creation of multiple choices from T-REx: TREx-MC
T-REx Elsahar et al. (2018) is a large-scale alignment dataset that aligns between Wikipedia abstracts and Wikipedia triples. We have utilized the processed version of T-REx available on HuggingFace 555https://huggingface.co/datasets/relbert/t_rex for our experiments. We filtered out the relations that have more than 500 facts and 100 unique object entities. The unique objects ensure having 100 feasible multiple-choices for each fact in each relation. We curated 50 relations for our dataset TREx-MC that essentially consists of subject, relation, multiple choices . The multiple choices comprise the correct answer along with 99 other potential choices. We list the 50 relations in Table 3 below.
@@ Appendix B Inference Setup
1.
Transformers Based Setup: This setup utilizes the utilities present in the transformers library Wolf et al. (2020) to obtain the log probabilities for generating the different options.
2.
@@
3.
SGLang Based Setup: SGLang Zheng et al. (2023) is a structured generation language designed for large language models (LLMs). It speeds up LLM interactions and provides enhanced control through tight integration of its frontend language and backend runtime system. SGLang also leverages Radix Attention to cache common components across queries in the KV cache, enabling substantial speedups. We observed sizable speed boosts for IC-LKE and EIC-LKE over vLLM. However, we are constrained by SGLang’s limited model family support at the moment, and only utilize it for the Llama, Mistral, and Mixtral families.
@@ Appendix C Implementation Details
C.1 IC-LKE
IC-LKE leverages 50 randomly chosen samples from the training data as in-context examples but does not use the relation name. The base prompt is now composed of 50 different examples followed by the name of the entity being tested. A sample would be “Albert Einstein 14 March 1879 Ernest Rutherford 30 August 1871 … J.J. Thomson 18 December 1856 Max Planck."
The subsequent process is the same as PB-LKE . The process involves adding 100 different choices to the base prompt. A single forward pass is conducted for each sequence, generating log probabilities for the entire sequence. The common part, represented by the tokens for the base prompt is then removed from the tokens of the concatenated base prompt and option resulting in the log probabilities for the option. Similar to PB-LKE , if the option is tokenized into multiple tokens, a single probability value is obtained by multiplying the individual token probabilities. The resulting values are normalized across multiple choices, and the option with the highest probability is selected as the correct answer. We use the vLLM Based & SGLang Based Setup for this LKE.
C.2 EIC-LKE
The EIC-LKE retrieves all 100 samples from our training dataset, initially maintaining them in a single sequence. Then, starting from the 50th training sample, we intersperse our test sample with all the choices every 5 examples. This results in a sequence that includes both the correct and incorrect choices. To determine the probability of each choice, we first use a tokenizer to tokenize all the subjects and choices separately. Then, we combine their token IDs, using a space token to separate the subject and object, and a comma to differentiate between different tuples. After obtaining the sequence’s token IDs, we input these token IDs into a simple forward pass. We use the token length of each subject and object to locate the probability of their corresponding tokens. Finally, we calculate the probability of all the choices by multiplying the probabilities of all their tokens. The resulting values are normalized across the choices, and the choice with the highest probability is selected as the correct answer. We use a vLLM Based Setup for this LKE.
@@ Appendix D Different Metrics
The evaluation metric can readily be adapted to existing classification metrics. For example, we introduced the metric Accuracy@K, a calibrated measure that assesses a model’s confidence in its predictions. This metric quantifies how accurately the model identifies knowledge at specified confidence levels for a given relation. We filter the instances that have their confidence levels > threshold and form the set . Following this, we use our accuracy measure to compute Accuracy@K for varying values of , the results of which are shown in Figure 8.
@@ Appendix E Probabilities of objects in sequence
We first consider correct examples (subject-object pairs) and report the absolute generation probability of objects in corresponding examples. We showed the results for Llama2-7B, Falcon-7B, Gemma-7B, and Pythia-12B in Figure 11, Figure 9 and Figure 10. Figure 11a, Figure 9a, and Figure 10a illustrates the probability of each object at various sequence positions; Figure 11b, Figure 9b, and Figure 10b shows the impact on probabilities after substituting 40 objects dispersed within the sequence with incorrect ones. Figure 11c, Figure 9c, and Figure 10c visualizes the effect of replacing objects at simultaneous positions. Figures 11d, Figure 9d, Figure 10d, Figure 11e, Figure 9e, and Figure 10e present the outcomes of using unknown subject-object pairs as replacements. We used a horizontal dashed line showing an average probability of the correct examples. The yellow star notated the example at position 114 in the sequence.
Appendix F Efficient In Context learning based LKE (EIC-LKE)
We improve the efficiency of IC-LKE to perform knowledge extraction of multiple test facts in a single prompt. Leveraging the context length in LLMs, the efficient version, namely EIC-LKE, places multiple test facts surrounded by training facts into the same prompt. We measure the object probability of each of the (alternative) test facts in the seuqence to determine whether the LLM assigns higher probability to the correct fact than the others.
Example 2.
Considering the training facts in Example 1, we evaluate two test choices (highlighted in yellow) for the birth-year relation: Einstein, birth-year 1879, and Louis birth-year, 1892, using two prompts instead of four as in IC-LKE.
| Feynman 1918 Heisenberg 1901 Louis 1850 | ||
| Feynman 1918 Einstein 1880 Heisenberg 1901 |
Appendix G Details about the human-generated prompts and machine-mined prompts
We list the used human-generated and machine-mined prompts from (Jiang et al., 2020) in Table 5 with subjects denoted as <‘head’>.
@@ Appendix G Details about the human-generated prompts and machine-mined prompts
| 1 | {subject} is signed to | {subject} signed with | |
| Record label | 2 | {subject} is a recording artist for | {subject} sohmed a recording contract with |
| 3 | {subject} is a recording artist on | {subject} released by | |
@@
Appendix H Additional results
H.1 Model Name Simplification
@@
H.2 Additional results on baseline comparison
We compare IC-LKE and EIC-LKE on 12 relations from T-REx-MC: capital, named after, developer, manufacturer, genre, instance of, native language, original broadcaster, language spoken written or signed, original language of film / TV show, official language, position played on team/speciality. We chose those 12 relations from T-REx-MCthat are found to be in common with Jiang et al. (2020) where they define the templates for HGP and MMP. We evaluated 4 models (Mistral-7B, Llama-7B, Falcon-7B, and Pythia-12B) and showed all the results in Figure 12.
H.3 Full order of models and relations
We evaluated 49 models on 50 relations by our IC-LKEand EIC-LKE. Table LABEL:table:model_order shows the ordered models by the average accuracy of all the 50 relations. Table 8 shows the ordered relations by the average accuracy of all the 49 models.
H.4 Full evaluation on EIC-LKE
We evaluated all the pre-trained models using EIC-LKE, but didn’t evaluate GPT-NEOX-20B due to the limitation of its context window size. Figure 13 shows the heatmap of models vs. relations, ordered in the same way as in Figure 5.
@@ H.4 Full evaluation on EIC-LKE
H.5 Relation accuracy correlation of all the pre-trained models
@@
H.6 Impact of finetuning
We show the results evaluated by EIC-LKE for all the pre-trained models and fine-tuned models in Figure 15 from the relations in TREx-MC, which also conveys the message about reduced knowledge in fine-tuned models.
We also show the results for the average subsumption rate () for base models and fine-tuned models over the relations in T-REx-MC.
@@ H.6 Impact of finetuning
H.7 Evaluation of Generated Output
We also evaluated the generated output, where we used greedy searching(temperature=0), and asked both pre-trained and fine-tuned models to generate 50 tokens using different prompts from HGP and MMP. Following this, we checked for the presence of the ground truth in the generated output of 50 tokens. The generation is correct if present, and incorrect otherwise, then we compute the generation accuracy on the test dataset. We report the average generation accuracy based on 12 relations and the HGP/MMP templates shown in Table 5.

### v2

@@ 4.2.2. Comparing within the same LLM family
Larger models embed more knowledge with certain exceptions. Figure 7 shows that within each model family, larger models (e.g., Llama-65B) generally outperform smaller ones (e.g., Llama-13B). Models within the same family are typically pre-trained on the same datasets (Biderman et al., 2023; Zhang et al., 2022; Touvron et al., 2023). The results suggest that, when trained on identical datasets, larger models in general capture a broader set of facts. An exception lies in the OPT group of models (Zhang et al., 2022) that have also been trained on the Pile dataset (Gao et al., 2020) like some of the other models (Biderman et al., 2023). This may call for investigating the ways of knowledge injection and validating if that can be attributed to such a performance deviation.
@@
Despite being trained on the same data, models might remember different facts. From the above results, it is not clear if the larger models are subsuming smaller models in their factual knowledge, i.e., do the larger models correctly identify the facts that the smaller models are correct on? To assess this, we compute the subsumption rate :
@@
that measures how much of the fraction of facts from known by smaller model are also recognised by the larger model . A subsumption rate of 1 indicates that all of the smaller model’s knowledge is also contained in the larger model.
@@
Table 2 shows the average subsumption rate () between the largest and smallest models in a family, as well as the average accuracy, over all relations for different model families. Interestingly, is relatively low (¡ 0.5) for OPT, Pythia and Bloom (i.e., the larger models know less than 50% of what the smaller models know) and only reaching up to 0.8 for Gemma, Llama and Llama-2. Therefore, even though models within each family are trained on the same datasets and generally agree on the relative knowledge of different relations (Figure 8), there are differences in the knowledge of specific facts they retain from their training data. These discrepancies suggest that simply increasing the model size may not be sufficient to enhance factual knowledge, thus requiring the need for proper factual knowledge injection into the models.
@@
Instruction fine-tuning reduces latent knowledge. Finally, we investigate the effects of chat-based instruction fine-tuning on the factual knowledge of models. Base language models are often fine-tuned (using a mix of supervised and reinforcement learning (Ouyang et al., 2022)) to improve their ability to follow instructions. While previous studies have shown that fine-tuning enhances performance on various benchmarks, its impact on latent knowledge is unclear. Figure 9 illustrates the comparative accuracy of pre-trained models and their fine-tuned counterparts. In almost all cases, the fine-tuned models obtain lower accuracy than their base versions, suggesting that fine-tuning reduces the amount of latent knowledge estimation. A similar observation was made by (Yu et al., 2024). To further assess if fine-tuned models acquire new knowledge, we compute the subsumption rate between pre-trained and fine-tuned versions (Table 11). We find that most latent knowledge in fine-tuned models is already present in base models (high ). This outcome highlights the need for caution when fine-tuning models, as these adjustments might inadvertently compromise with the existing internal knowledge.
5. Concluding Discussion
@@
In this work, we investigate a new way to estimate latent factual knowledge from an LLM. Unlike prior approaches, our method does not engineer prompts (zero-prompting). Rather it relies on LLMs’ in-context learning ability to infer the factual knowledge question and the expected answer format. Our method not only addresses many reliability concerns with prompting, but it also recollects significantly more factual knowledge than prompting. In contrast to prompting, which requires relationship-specific and LLM-specific prompt engineering, Our method can be applied with minimal effort to test factual knowledge of relations across a variety of structured knowledge bases and LLMs. This ability enables us to compare the latent knowledge captured by many different families of open-source LLMs; we expect our results to be of interest to the designers of these LLMs. Finally, to design our zero-prompt many-shot LKE, we explore the impact of the number and order of correct, incorrect, and unknown examples used as inputs; our findings may be of independent interest to developing a better understanding of different learning modes of in-context learning.
@@
A fundamental question posed by our and prior work on estimating latent knowledge in LLMs: What does it mean for an LLM to know a fact? Suppose we tried to infer if an LLM knows the capital of Germany using the input ”France Paris; Spain Madrid; Germany ” and suppose the answer was Berlin. What we have learned is that the LLM knows that the relationship between Germany and Berlin is similar to that between France and Paris or Spain and Madrid. What we have not learned is whether the LLM knows that the relation is called ”capital” in English or ”hauptstadt” in German. The latter is revealed by prompts such as ”The capital of Germany is ”. But, such prompts don’t reveal whether the LLM knows that what Berlin means to Germany is similar to what Paris means to France.
@@
Is one type of knowing facts better than another? It is difficult to answer in general. Neither type of knowing guarantees that the knowledge can be put to use in different contexts and tasks, such as when we ask the LLM where the parliament of Germany is located. However, they lead to different strategies for getting LLMs to generate correct outputs. With the first type of knowing, we can use a list of facts as input such as ”The parliament of France is in Paris; The parliament of Spain is in Madrid; The parliament of Germany is in ”. With the second type of knowing, we can hope to use a chain of thought prompts such as ”The parliament of a country is in its capital. The parliament of Germany is in ”. Nevertheless, one clear takeaway from our study is related to how factual knowledge is latently embedded in an LLM. We show that more factual knowledge can be recollected using in-context learning, i.e., the representations of subjects and objects that share the same relationship, than by prompting with the name of their relationship.
6. Ethical Considerations
Our research utilizes public datasets and open-source LLMs, which mitigates immediate privacy concerns. However, our findings on the factual knowledge capabilities of various LLMs could influence their deployment in real-world applications, potentially leading to over-reliance on models for tasks requiring factual accuracy. We encourage users of our methodology to consider these implications and to use the knowledge estimation techniques responsibly, with appropriate safeguards against potential misuse. Furthermore, as our work may reveal biases or gaps in the factual knowledge of LLMs, we urge developers to address these issues to ensure fair and equitable AI systems.
@@ Appendix A Dataset
A.1. Creation of Nobel laureates dataset from Wikidata
@@
A.2. Creation of multiple choices from T-REx: TREx-MC
T-REx (Elsahar et al., 2018) is a large-scale alignment dataset that aligns between Wikipedia abstracts and Wikipedia triples. We have utilized the processed version of T-REx available on HuggingFace 555https://huggingface.co/datasets/relbert/t_rex for our experiments. We filtered out the relations that have more than 500 facts and 100 unique object entities. The unique objects ensure having 100 feasible multiple choices for each fact in each relation. We also manually filtered out relations with multiple correct objects (e.g. “America”, “USA”, “American”) to avoid ambiguity. Additionally for relations that have objects in the form of partial matches (e.g. “French”, “French language”), the respective objects have been standardized to uniform values (e.g. “French”). We curated 50 relations for our dataset TREx-MC that essentially consists of subject, relation, multiple choices . The multiple choices comprise the correct answer along with 99 other potential choices. We list the 50 relations in Table 4 below.
@@ Appendix B Inference Setup
Transformers Based Setup: This setup utilizes the utilities present in the transformers library (Wolf et al., 2020) to obtain the log probabilities for generating the different options.
@@
SGLang Based Setup: SGLang (Zheng et al., 2023) is a structured generation language designed for large language models (LLMs). It speeds up LLM interactions and provides enhanced control through tight integration of its frontend language and backend runtime system. SGLang also leverages Radix Attention to cache common components across queries in the KV cache, enabling substantial speedups. We observed sizable speed boosts for ZP-LKE over vLLM. However, we are constrained by SGLang’s limited model family support at the moment, and only utilize it for the Llama, Mistral, and Mixtral families.
@@ Appendix C Implementation Details
ZP-LKE leverages 50 randomly chosen samples from the training data as in-context examples but does not use the relation name. The base prompt is now composed of 50 different examples followed by the name of the entity being tested. A sample would be “Albert Einstein 14 March 1879 Ernest Rutherford 30 August 1871 … J.J. Thomson 18 December 1856 Max Planck.”
A single forward pass is conducted for each sequence, generating log probabilities for the entire sequence. The common part, represented by the tokens for the base prompt is then removed from the tokens of the concatenated base prompt and option resulting in the log probabilities for the option. If the option is tokenized into multiple tokens, a single probability value is obtained by multiplying the individual token probabilities. The resulting values are normalized across multiple choices, and the option with the highest probability is selected as the correct answer. We use the vLLM Based & SGLang Based Setup for this LKE.
@@ Appendix D Different Metrics
The evaluation metric can readily be adapted to existing classification metrics. For example, we introduced the metric Accuracy@K, a calibrated measure that assesses a model’s confidence in its predictions. This metric quantifies how accurately the model identifies knowledge at specified confidence levels for a given relation. We filter the instances that have their confidence levels ¿ threshold and form the set . Following this, we use our accuracy measure to compute Accuracy@K for varying values of , the results of which are shown in Figure 10.
@@ Appendix E Probabilities of objects in sequence
We first consider correct examples (subject-object pairs) and report the absolute generation probability of objects in corresponding examples. We showed the results for Llama2-7B, Falcon-7B, and Pythia-12B in Figure 11, Figure 12 and Figure 13. Figure 11(a), Figure 12(a) and Figure 13(a) illustrates the probability of each object at various sequence positions; Figure 11(b), Figure 12(b) and Figure 13(b) shows the impact on probabilities after substituting 40 objects dispersed within the sequence with incorrect ones. Figure 11(c), Figure 12(c) and Figure 13(c) visualizes the effect of replacing objects at simultaneous positions. Figure 11(d), Figure 12(d), Figures 13(d), Figure 11(e), Figure 12(e) and Figure 13(e) present the outcomes of using unknown subject-object pairs as replacements. We used a horizontal dashed line showing an average probability of the correct examples. The yellow star marks the example at position 114 in the sequence.
Appendix F Details about the human-generated prompts and machine-mined prompts
We list the used human-generated and machine-mined prompts from (Jiang et al., 2020) in Table 6 with subjects denoted as ¡‘subject’¿.
@@ Appendix F Details about the human-generated prompts and machine-mined prompts
| 1 | in order to communicate in | ||
| Language spoken, written or signed | 2 | {subject} used to communicate in | |
| 3 | |||
@@
Appendix G Additional results
G.1. Model Name Simplification
@@
G.2. Additional results on baseline comparison
We compare ZP-LKE on 12 relations from T-REx-MC: capital, named after, developer, manufacturer, genre, instance of, native language, original broadcaster, language spoken written or signed, original language of film / TV show, official language, position played on team/speciality. We chose those 12 relations from T-REx-MC that are found to be in common with (Jiang et al., 2020) where they define the templates for HGP and MMP. We evaluated 4 models (Mistral-7B, Llama-7B, Falcon-7B, and Pythia-12B) and showed all the results in Figure 14 and Figure 15.
G.3. Full order of models and relations
We evaluated 49 models on 50 relations by our ZP-LKE . Table LABEL:table:model_order shows the ordered models by the average accuracy of all the 50 relations. Table 9 shows the ordered relations by the average accuracy of all the 49 models.
@@ G.3. Full order of models and relations
G.4. Relation accuracy correlation of all the pre-trained models
@@
G.5. Impact of finetuning
We also show the results for the average subsumption rate () in Table 11 for base models and fine-tuned models over the relations in T-REx-MC.